OpenStudy (anonymous):

3^x = (1/9)^3 What is the value of x?

OpenStudy (anonymous):

\[3^x=(\frac19)^3\] \[3^x=(9^{-1})^3=9^{-3}=(3^2)^{-3}=3^{-6}\] x=-6

OpenStudy (anonymous):

How did you get the 3^2?